OpenStudy (anonymous):

Use the function to answer the question. f(x) = 3|x + 5| + 8 What is the vertex of the graph of f(x)?

OpenStudy (anonymous):

I think its (5,8)

OpenStudy (anonymous):

solve y=3(x+5) +8 and y=-3(x+5) +8

OpenStudy (anonymous):

(-5,8)

OpenStudy (anonymous):

Ok @akash123 why is it -5?

OpenStudy (anonymous):

solve y=3(x+5) +8 and y=-3(x+5) +8 u'll get (-5,8)

OpenStudy (anonymous):

use I x+5 I = -(x+5), x<-5 =(x+5), x >-5